OpenStudy (anonymous):

If possible, completely factor the expression. (If the polynomial is not factorable using integers, enter PRIME.)

OpenStudy (ganpat):

expression ?

OpenStudy (anonymous):

sry 2^2+13x+40

OpenStudy (ganpat):

is it 2^2 or x^2 ??

OpenStudy (anonymous):

its x squared

OpenStudy (anonymous):

i am sorry not typing very well today

OpenStudy (ganpat):

no probs.. x2+13x+40 = 0 x2 + 8x + 5x + 40 = 0 as, 8 * 5 = 40 .. and 8 + 5 = 13. can u solve further ??

OpenStudy (anonymous):

yes thank you

OpenStudy (ganpat):

cool !!!

OpenStudy (ganpat):

x2 + 8x + 5x + 40 = 0 x (x +8) + 5 (x +8) = 0 (x +5) (x +8) = 0 x = -5 and x = -8...